OpenStudy (anonymous):

what is the area,in square units,of a square that has the same perimeter as the rectangle above Width=5 length=9 i dont know what is the question asking

OpenStudy (anonymous):

perimeter of a rectangle = 2(l+b) = 2(5+9) =2(14) = 28 perimeter of rectangle = perimeter of square 28 = 4s 4s = 28 s = 28/4 s = 7 area of square=s^2 = 7*7 = 49

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

mathematicalpro is right!

OpenStudy (anonymous):

thanks for confirming my answer jurasarts